OpenStudy (anonymous):

5 men are hired to complete a job. If an additional man is hired, the job can be completed 8 days earlier. Given that the number of days required to complete the job is inversely proportional to the number of men hired, find the number of additional men who must be hired in order for the job to be completed 28 days earlier.

OpenStudy (anonymous):

5 men => n/5 days 6men => n/6 days n/5 - n/6 = n/30 days = 8days n = 240 by 5 men it will be finish it 48 days 28 days earlier = 20 days 240 /20 = 12 12 - 5 = 7 additional men